Calculate the present value of the compound interest loan. (Round your answers to the nearest cent.)

$26,000 after 6 years at 5% if the interest is compounded in the following ways.

(a) annually
$  

(b) quarterly

$
Answer #1
PV = FV/(1+r/m)^(m*n)
where r = the annual interest
m = the frequency of compounding in a year
n = number of years.
a) PV = 26000/1.05^6 = $   19,401.60
b) PV = 26000/(1+0.05/4)^(6*4) = $   19,297.12